17. Find the slope of the line that is (a) parallel and (b) perpendicular to the line through the pair of points.
(-7,-2) and (1,7)
Type an integer or simplified fraction.

Answer by mananth(16946)

(-7,-2) and (1,7)
...
slope m of a line = (y2-y1)/(x2-x1)
m=(7-(-2))/(1-(-7))
m=9/8
..
The line parallel to the line passing through above points will have same slope
m= 9/8
...
For line perpedicular it will have a slope which is negative reciprocal of the slope
m=-8/9
...
m.ananth@hotmail.ca